OpenStudy (anonymous):

Find the lateral area of a right rectangular prism with width 5, length 6, and height 9.

hero (hero):

LA = 2(ab + bc + ac) LA = 2(30 + 54 + 45) LA = 2(129) LA = 258

OpenStudy (anonymous):

what is ab bc and ac mean?

hero (hero):

a = width b = length c = height

OpenStudy (anonymous):

Got it! Thanks!

OpenStudy (anonymous):

would that be the same thing as surface area then?